OpenStudy (anonymous):

Find an equation of the line. Horizontal; through (-3,-11) Write the equation in standard form.

OpenStudy (anonymous):

This is vectors right?

OpenStudy (anonymous):

what is vectors ?

OpenStudy (anonymous):

...calculus, okay i can still do this using simper math are you familiar with this equation = y-y1 =m(x-x1)

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

Okay, you have y1, and x1 already from the point given (-3,-11) and m is slope, what is the slope of a horizontal line?

OpenStudy (anonymous):

the slope is 0 ?

OpenStudy (anonymous):

yup so y-(-11)=0 (because anything multiplied by zero is zero and then basically you have y + 11=0 so y=-11

OpenStudy (anonymous):

in standard form though its just y+11=0

OpenStudy (anonymous):

oh thank you

OpenStudy (anonymous):

no problem